OpenStudy (anonymous):

2500cu.cm of hydrogen at STP is subjected to a 2-1/2 times increase in pressure, temp remaining constant. What volume will the hydrogen now occupy? As per Charles' law the answer is 1000cu.cm but the answer key shows 5000/7 cu.cm. Which is correct?

OpenStudy (matt101):

You want to use Boyle's Law, P1V1 = P2V2, rather than Charles' Law (which relates volume and temperature) since all variables stay constant after the change with the exception of pressure and volume. The final pressure is 2.5 times greater than the initial pressure, i.e. P2 = 2.5 P1. Therefore: P1V1 = P2V2 P1V1 = 2.5P1V2 V1 = 2.5V2 V2 = (V1)/2.5 = 2500/2.5 = 1000 cu.cm You'll notice that this equation is not dependent on the units of either pressure or volume, and that pressure and volume are inversely proportional to one another (that is if one increases by a factor of 2.5, the other decreases by a factor of 2.5, for example). Your answer is right, the book answer is weird.

OpenStudy (anonymous):

Thanks Matt...